# Energy exchange between electrons and ions driven by ITG-TEM turbulence

### v1

@@ I Introduction
Turbulence in magnetically confined plasmas induces particle flux, heat flux, and energy exchange between electrons and ionsKato2024 ; Sugama1996 ; Sugama2009 ; Waltz ; Waltz1997 ; Candy ; Manheimer . Gyrokinetic theory provides a robust framework for understanding turbulence in magnetized plasmas, and gyrokinetic simulations are widely used to analyse the turbulent transportsAntonsen ; CTB ; F-C ; Pueschel ; Dimits ; Krommes2012 ; Garbet2010 ; Idomura2006 ; Horton . The prediction of turbulent effects is necessary for simulations of global density and temperature profiles. The quasilinear model is useful for estimating the nonlinear simulation results by linear simulations, which require relatively low computational costsCasati ; Citrin ; Citrin2017 ; Bourdelle ; Staebler2021 ; Staebler2024 ; Merz ; Narita ; Toda ; Parker ; Barnes .
It is well established that the turbulent fluxes significantly contribute to the global profiles. However, the effect of turbulent energy exchange had been considered negligible Waltz ; Waltz1997 ; Candy ; Manheimer . As a result, while turbulent transport fluxes are commonly incorporated into transport simulations for predicting density and temperature profiles, the contribution of turbulent energy exchange is typically omitted.
Our previous work Kato2024 , which investigates energy exchange caused by pure ITG turbulence in tokamak plasmas, has reported that in high temperature or low collisional plasmas, such as those handled in future fusion reactors, turbulent energy exchange could exceed collisional one. Furthermore, it is found that turbulent energy exchange in ITG turbulence can transfer energy from ions to electrons regardless of whether electrons or ions are hotter. This is significantly different from the collisional energy exchange, which transfers energy from the hotter particle species to the cooler species. This implies that the ITG turbulence should be suppressed from the viewpoint of sustaining the high ion temperature required for fusion reactions since it prevents energy transfer from alpha-heated electrons to ions. Thus, transport simulations predicting global temperature profiles should include the effect of turbulent energy exchange for future fusion reactors. To support this, the feasibility of the quasilinear model for energy exchange due to ITG turbulence in tokamak plasmas has also been demonstrated.
@@
The rest of this paper is organized as follows. In Sec. II, the entropy balance equation for the turbulent fluctuation and components of turbulent energy transfer are introduced. In Sec. III, simulation results of pure TEM and mixed ITG-TEM turbulence by the GKV codeGKV , which uses a flux tube domainBeer , are shown. Simulation settings and results of linear simulations are shown in Sec. III.1. The nonlinear simulation results for pure TEM and mixed ITG-TEM turbulence are presented in Secs. III.2 and III.3, respectively, and the entropy balance, energy transfer and the feasibility of the quasilinear model for energy transfer are discussed. Finally, conclusions and discussion are given in Sec. IV.
II Entropy balance equation and turbulent energy transfers
We here introduce the balance equation for the entropy associated with the perturbed distribution function and the components of turbulent energy transfers for electrons and ions. In this study, the background radial electric field $E_{r}$ is set to zero, since its effect appears only in the Doppler shift in the local gyrokinetic theory and does not fundamentally affect the turbulent transport fluxes and energy exchange. The background $E\times B$ flow shear effect is also neglected here since it is regarded as smaller by an order of the gyrokinetic expansion parameter $\delta$ than the terms included in the gyrokinetic equation introduced in Ref. Kato2024 . The detailed derivation is explained in Refs. Kato2024 ; Sugama1996 ; Sugama2009 .
The entropy balance equation in the perpendicular wavenumber space for particle species $a=e,i$ is described as,
@@ II Entropy balance equation and turbulent energy transfers
where the perturbed entropy for nonadiabatic distrbution function $\delta S_{ha\bm{k}}$, particle flux $\Gamma_{a\bm{k}}^{\rm turb}$, heat flux $q^{\rm turb}_{a\bm{k}}$, energy transfer from the perturbed field to particles $Q^{\rm turb}_{a\bm{k}}$, collisional dissipation $D_{a\bm{k}}$, and entropy transfer by nonlinear interaction $N_{a\bm{k}}$, pressure and temperature gradient scale lengths $L_{pa}$, $L_{Ta}$ are described as,
@@
×$\hskip 71.13188pt\times h_{a\bm{k}}^{*}\left(-i\frac{c}{B}\psi_{a\bm{k}}\bm{k}\times\bm{b}\right)\cdot\nabla r\Bigg{\rangle}\Bigg{\rangle},$ (3)
@@
respectively. The superscript * and "$\rm Re$" denote the complex conjugate and real parts, respectively, and $c$, $e_{a}$, $m_{a}$, $T_{a}$, $p_{a}$, $f_{Ma}$, and $C^{GK}_{a}$ are the speed of light, charge, mass, temperature, pressure, local Maxwellian, and gyrokinetic collision operator, respectively. The perturbed nonadiabatic distribution function with a perpendicular wavenumber $\bm{k}$, denoted as $h_{a\bm{k}}$, is related to the perturbed distribution function $f_{a\bm{k}}$ by the following relation:
[…]
$\psi_{a\bm{k}}=J_{0}\left(k\rho_{a}\right)\left(\phi_{\bm{k}}-\frac{v_{\parallel}}{c}A_{\parallel\bm{k}}\right)+J_{1}\left(k\rho_{a}\right)\frac{v_{\perp}}{c}\frac{B_{\parallel\bm{k}}}{k},\$
@@
where $\sum_{\bm{k}}N_{a\bm{k}}=0$. Under the transport time scaling $\partial\langle\langle\cdots\rangle\rangle/\partial t=\mathcal{O}(\delta^{2})$, the L.H.S. of Eq. (9) is negligible. Then, the four terms in the R.H.S of Eq. (9) should be balanced. In addition, the sum of energy transfer over particle species $\sum_{a=e,i}Q^{\rm turb}_{a}$ also becomes negligible under the transport time scaling. Therefore, the energy transfer from the turbulent fluctuation to particle species $a$, $Q^{\rm turb}_{a}=\sum_{\bm{k}}Q^{\rm turb}_{a\bm{k}}$, can be regarded as the energy exchange between different particle species.
According to Ref.Kato2024 , the direction of turbulent energy transfer, ${Q^{\rm turb}_{a}}$, in ITG turbulence is not determined by the temperature difference between species, in contrast to collisional energy exchange. In other words, ITG turbulence can transfer energy from cooler ions to hotter electrons, which appears to contradict the second law of thermodynamics. However, this seemingly paradoxical phenomenon is explained by the fact that the entropy production associated with ion particle and heat fluxes exceeds that of the electrons, thereby demonstrating that it does not contradict the second law. Based on this observation, it is conjectured that energy is generally transferred by turbulence from a particle species with larger entropy production due to particle and heat transport driven by the instability to other species, regardless of which species is hotter. This study examines the validity of this conjecture in the context of ITG–TEM turbulence.
The turbulent energy transfer in the wavenumber space $Q^{\rm turb}_{a\bm{k}}$ can be divided into four components under the transport time scaling, described as
@@
The collisional components $Q^{\rm turb}_{aC\bm{k}}(a=e,i)$ are negligible in low collisional plasmas, and the component of nonlinear interaction $Q^{\rm turb}_{a\psi\bm{k}}$ satisfies
@@
Then, the turbulent energy transfer $Q^{\rm turb}_{a}$ is determined by the Joule heating via currents resulting from parallel streaming motions to the background magnetic field, $Q^{\rm turb}_{a\parallel}=\sum_{\bm{k}}Q^{\rm turb}_{a\parallel\bm{k}}$, and from $\nabla B$-curvature drifts perpendicular to it, $Q^{\rm turb}_{aB}=\sum_{\bm{k}}Q^{\rm turb}_{aB\bm{k}}$.
@@
Charge ratio $e_{e}/e_{i}$1
Ion beta value $\beta_{i}=4\pi n_{i}T_{i}/B^{2}$ $1\times 10^{-4}$
@@
Collision frequency for electrons $\nu^{*}_{ee}\equiv R_{0}q_{0}\nu_{ee}/(\epsilon_{r}^{3/2}v_{ti})$ 0.007
@@ III.1 Simulation settings
In this study, microturbulence simulations are conducted using the GKV code GKV , which numerically solves the gyrokinetic equations within the flux-tube domain. Table LABEL:tab:Plasma_and_field_parameters shows plasma and field parameters, and Tab. LABEL:tab:resolution shows resolution settings used in this work. We here investigate ITG-TEM turbulence, and set the temperature ratio $T_{e}/T_{i}=3.0$ to suppress the ETG instability Nakata . The normalized electron and ion temperature gradients, $R_{0}/L_{Te}$ and $R_{0}/L_{Ti}$, are varied in order to modify the dominant instability while keeping the total pressure gradient, $\sum_{a=e,i}(R_{0}/L_{na}+R_{0}/L_{Ta})$, constant. To focus on pure TEM turbulence and suppress instabilities in high wavenumber modes, we set $(R_{0}/L_{Te},R_{0}/L_{Ti})=(7.0,1.0)$ instead of using $R_{0}/L_{Ti}=0.0$. We set the normalized density gradients $R_{0}/L_{na}=2.22(a=e,i)$, which is the same parameter used in the CBC caseDimits . The flux tube coordinates used here are $x=r-r_{0},y=r_{0}/q_{0}\left(q(r)\theta-\zeta\right)$, and $z=\theta$ in low-$\beta$, large aspect ratio, axisymmetric tokamak plasmas with circular, concentric flux surfaces, where $r$, $\theta$, and $\zeta$ denote the minor radius, poloidal angle, and toroidal angle, respectively. The subscript 00 signifies the parameters in the center of the flux tube.
@@
The perpendicular wavenumber is expressed as $\bm{k}=\left(k_{x}+\hat{s}zk_{y}\right)\bm{e}_{r}+k_{y}\bm{e}_{\theta}$. Here, $k_{x}$ and $k_{y}$ represent wavenumbers along the directions of $\nabla x$ and $\nabla y$, respectively, while $\bm{e}_{r}$ and $\bm{e}_{\theta}$ are unit vectors aligned with $\nabla r$ and $\nabla\theta$, respectively. The ion beta value is set to $\beta_{i}=4\pi n_{i}T_{i}/B^{2}=1\times 10^{-4}$, ensuring the validity of the electrostatic approximation. In these simulations, $B_{\parallel\bm{k}}$ is neglected, whereas $A_{\parallel\bm{k}}$ is retained to prevent numerical difficulties arising from very rapid electrostatic waves known as the $\omega_{H}$ mode Lee . The Lenard-Bernstein collision operator Lenard is employed due to its computational efficiency compared to more sophisticated collision models. Nevertheless, we anticipate that the choice of collision model does not significantly affect the results of this study, given that the normalized collision frequency is set to $\nu^{*}_{ii}\equiv R_{0}q\nu_{ii}/(\epsilon^{3/2}v_{ti})=0.068$, which is considerably smaller than the growth rates of the instabilities examined here.
@@
Fig. 1 shows linear growth rates and real frequencies obtained by linear simulations varying the temperature gradients. The sign of real frequencies reverses around $(R_{0}/L_{Te},R_{0}/L_{Ti})=(5.0,3.0)$, indicating that the most unstable mode transitions between ITG and TEM. Nonlinear simulations are performed using these plasma parameters to investigate the direction, magnitude, underlying physical mechanisms, and predictability of quasilinear models for turbulent energy exchange in ITG-TEM turbulence.

### v2

@@ I Introduction
Turbulence in magnetically confined plasmas induces particle flux, heat flux, and energy exchange between electrons and ionsKato2024 ; Sugama1996 ; Sugama2009 ; Waltz ; Waltz1997 ; Candy ; Manheimer . Gyrokinetic theory provides a robust framework for understanding turbulence in magnetized plasmas, and gyrokinetic simulations are widely used to analyse the turbulent transportAntonsen ; CTB ; F-C ; Pueschel ; Dimits ; Krommes2012 ; Garbet2010 ; Idomura2006 ; Horton . The prediction of turbulent effects is necessary for simulations of global density and temperature profiles. The quasilinear model is useful for estimating the nonlinear simulation results by linear simulations, which require relatively low computational costsCasati ; Citrin ; Citrin2017 ; Bourdelle ; Staebler2021 ; Staebler2024 ; Merz ; Narita ; Toda ; Parker ; Barnes .
It is well-established that the global profiles are predominantly determined by a balance between sources of particles and energy, and losses due to turbulent fluxes. However, the effect of turbulent energy exchange had been considered negligible Waltz ; Waltz1997 ; Candy ; Manheimer . As a result, while turbulent transport fluxes are commonly incorporated into transport simulations for predicting density and temperature profiles, the contribution of turbulent energy exchange is typically omitted.
Our previous work Kato2024 , which investigates energy exchange caused by pure ITG turbulence in tokamak plasmas, has reported that in high temperature or weakly collisional plasmas, such as those expected in future fusion reactors, turbulent energy exchange could exceed collisional one. Furthermore, it is found that turbulent energy exchange in ITG turbulence can transfer energy from ions to electrons regardless of whether electrons or ions are hotter. This is significantly different from the collisional energy exchange, which transfers energy from the hotter particle species to the cooler species. In other words, ITG turbulence can transfer energy from cooler ions to hotter electrons, which appears to contradict the second law of thermodynamics. However, this seemingly paradoxical phenomenon is explained by the fact that the entropy production associated with ion particle and heat fluxes exceeds that of the electrons, thereby demonstrating that it does not contradict the second law. Based on this observation, it is conjectured that energy is generally transferred by turbulence from a particle species with larger entropy production due to particle and heat transport driven by the instability to other species, regardless of which species is hotter. This study examines the validity of this conjecture in the context of ITG–TEM turbulence.
Since ITG turbulence not only enhances particle and heat transport but also inhibits the energy transfer from alpha-heated electrons to ions, suppression of ITG turbulence is favorable for sustaining the high ion temperature required for fusion reactions. Thus, transport simulations predicting global temperature profiles should include the effect of turbulent energy exchange for future fusion reactors. To support this, the feasibility of the quasilinear model for energy exchange due to ITG turbulence in tokamak plasmas has also been demonstrated in our previous work.
@@
The rest of this paper is organized as follows. In Sec. II, the entropy balance equation for the turbulent fluctuation and components of turbulent energy transfer are introduced. Following the previous paperKato2024 , the symbol $Q$ is used to denote the energy exchange while the energy flux is denoted by $\mathcal{E}$ in this paper. In Sec. III, simulation results of pure TEM and mixed ITG-TEM turbulence by the GKV codeGKV , which uses a flux tube domainBeer , are shown. Simulation settings and results of linear simulations are shown in Sec. III.1. The nonlinear simulation results for pure TEM and mixed ITG-TEM turbulence are presented in Secs. III.2 and III.3, respectively, and the entropy balance and energy transfer are discussed. The feasibility of the quasilinear model for energy transfer in ITG-TEM turbulence is investigated in Sec. III.4. Finally, conclusions and discussion are given in Sec. IV.
II Theoretical model
II.1 Entropy balance equation
We here introduce the balance equation for the entropy associated with the perturbed distribution function.
The perturbed nonadiabatic distribution function with a perpendicular wavenumber $\bm{k}$ for particle species $a=e,i$, denoted as $h_{a\bm{k}}$, is related to the perturbed distribution function $f_{a\bm{k}}$ by the following relation:
where $\bm{\rho}_{a}=\bm{b}\times\bm{v}/\Omega_{a}$, $\Omega_{a}=e_{a}B/(m_{a}c)$, $\bm{b}=\bm{B}/B$, $B=|\bm{B}|$ and $c$, $e_{a}$, $m_{a}$, $T_{a}$, $f_{Ma}$, and $\bm{B}$ are the speed of light, charge, mass, temperature, local Maxwellian, and background magnetic field, respectively.
$\psi_{a\bm{k}}=J_{0}\left(k\rho_{a}\right)\left(\phi_{\bm{k}}-\frac{v_{\parallel}}{c}A_{\parallel\bm{k}}\right)+J_{1}\left(k\rho_{a}\right)\frac{v_{\perp}}{c}\frac{B_{\parallel\bm{k}}}{k},\\$ (2)
In this study, the background radial electric field $E_{r}$ is set to zero, since its effect appears only in the Doppler shift in the local gyrokinetic theory and does not fundamentally affect the turbulent transport fluxes and energy exchange. The background $E\times B$ flow shear effect is also neglected here since it is regarded as smaller by an order of the gyrokinetic expansion parameter $\delta$ than the terms included in the gyrokinetic equation introduced in Ref. Kato2024 . The detailed derivation is explained in Refs. Kato2024 ; Sugama1996 ; Sugama2009 .
The entropy balance equation in the perpendicular wavenumber space is described as,
@@ II.1 Entropy balance equation
where the perturbed entropy for nonadiabatic distribution function $\delta S_{ha\bm{k}}$, particle flux $\Gamma_{a\bm{k}}^{\rm turb}$, heat flux $q^{\rm turb}_{a\bm{k}}$, energy transfer from the perturbed field to particles $Q^{\rm turb}_{a\bm{k}}$, collisional dissipation $D_{a\bm{k}}$, and entropy transfer by nonlinear interaction $N_{a\bm{k}}$, pressure and temperature gradient scale lengths $L_{pa}$, $L_{Ta}$ are described as,
@@
$\hskip 71.13188pt\times h_{a\bm{k}}^{*}\left(-i\frac{c}{B}\psi_{a\bm{k}}\bm{k}\times\bm{b}\right)\cdot\nabla r\Bigg\rangle\Bigg\rangle,$ (5)
@@
respectively. The superscript * and $\rm Re$" denote the complex conjugate and real parts, respectively, and $p_{a}$ and $C^{GK}_{a}$ are the pressure and gyrokinetic collision operator, respectively. The double angle bracket $\langle\langle\cdots\rangle\rangle$ denotes the double average over the ensemble and flux surface, and $v$ is the absolute value of the velocity $\bm{v}$. The minor radius $r$ is used as a label of flux surfaces of a toroidal plasma.
@@
where $\sum_{\bm{k}}N_{a\bm{k}}=0$. Under the transport time-scale orderingSugama1996 ; Sugama2009 ; Hinton1976 $\partial\langle\langle\cdots\rangle\rangle/\partial t=\mathcal{O}(\delta^{2})$, the L.H.S. of Eq. (10) is negligible. Then, the four terms in the R.H.S of Eq. (10) should be balanced. In addition, using Poisson’s equation and Ampere’s law in the gyrokinetic formKato2024 ; Sugama1996 ; Sugama2009 , the sum of energy transfer over particle species is written as
$\sum_{a}{Q^{\rm turb}_{a}}=\frac{1}{8\pi}\frac{\partial}{\partial t}\sum_{\bm{k}}\Bigg\langle\Bigg\langle\left(k^{2}+\lambda_{D}^{-2}\right)|\phi_{\bm{k}}|^{2}-|\bm{B}_{\bm{k}}|^{2}\Bigg\rangle\Bigg\rangle,\\$ (11)
where $\bm{B}_{\bm{k}}=i\bm{k}\times\bm{A}_{\bm{k}}$, and $\lambda_{D}=1/\sqrt{\sum_{a}{4\pi n_{a}e_{a}^{2}/T_{a}}}$ is the Debye length. The R.H.S. of Eq. (11) also becomes negligible under the transport time-scale ordering. Therefore, the energy transfer from the turbulent fluctuation to particle species $a$, $Q^{\rm turb}_{a}=\sum_{\bm{k}}Q^{\rm turb}_{a\bm{k}}$, can be regarded as the energy exchange between different particle species.
II.2 Energy flux and energy exchange driven by microturbulence
The turbulent energy transfer in the wavenumber space $Q^{\rm turb}_{a\bm{k}}$ can be divided into four components under the transport time-scale ordering, described as
@@ II.2 Energy flux and energy exchange driven by microturbulence
The collisional components $Q^{\rm turb}_{aC\bm{k}}(a=e,i)$ are negligible in weakly collisional plasmas, and the component of nonlinear interaction $Q^{\rm turb}_{a\psi\bm{k}}$ satisfies
@@
Then, the turbulent energy transfer $Q^{\rm turb}_{a}$ is determined by the Joule heating via currents resulting from streaming motions parallel to the background magnetic field, $Q^{\rm turb}_{a\parallel}=\sum_{\bm{k}}Q^{\rm turb}_{a\parallel\bm{k}}$, and from $\nabla B$-curvature drifts perpendicular to it, $Q^{\rm turb}_{aB}=\sum_{\bm{k}}Q^{\rm turb}_{aB\bm{k}}$.
@@
II.3 Quasilinear model
The quasilinear model estimates the nonlinear simulation results based on two components: (i) the ratio of fluxes or energy exchange to the squared amplitude of the electrostatic potential, obtained from linear simulations, and (ii) the saturation rule of the squared amplitude of the electrostatic potential in a steady state turbulence. In this study, we focus on the former and do not deal with the latter. The ratio is calculated by
where $X_{a}$ indicates $(\Gamma_{a},q_{a},Y_{a})$. The superscript 0 and subscript $L,N$ denote the results obtained using only the $k_{x}=0$ modes (instead of summing over the $k_{x}$ space), those from linear and nonlinear calculations, respectively. Normalization factors $C_{X}$, and the real parts of integrals inside the double average over the ensemble and the flux surface $\langle\langle\cdots\rangle\rangle$ of Eqs. (5), $\tilde{X}_{a\bm{k}}=(\tilde{\Gamma}_{a\bm{k}},\tilde{q}_{a\bm{k}},\tilde{Y}_{a\bm{k}})$, are described as
@@ II.3 Quasilinear model
Charge ratio $e_{e}/e_{i}$ -1
Ion beta value $\beta_{i}=4\pi n_{i}T_{i}/B_{ax}^{2}$ $1\times 10^{-4}$
@@
Collision frequency for electrons $\nu^{*}_{ee}\equiv R_{0}q_{0}\nu_{ee}/(\epsilon_{r}^{3/2}v_{te})$ 0.007
@@ III.1 Simulation settings
In this study, microturbulence simulations are conducted using the GKV code GKV , which numerically solves the gyrokinetic equations within the flux-tube domain. Table LABEL:tab:Plasma_and_field_parameters shows plasma and field parameters, and Tab. LABEL:tab:resolution shows resolution settings used in this work. We here investigate ITG-TEM turbulence, and set the temperature ratio $T_{e}/T_{i}=3.0$ to suppress the ETG instability Nakata . The normalized electron and ion temperature gradients, $R_{0}/L_{Te}$ and $R_{0}/L_{Ti}$, are varied in order to modify the dominant instability while keeping the total pressure gradient, $\sum_{a=e,i}(R_{0}/L_{na}+R_{0}/L_{Ta})$, constant. The pressure-gradient component of the drift velocity is not neglected but remains constant in this study. To focus on pure TEM turbulence and suppress instabilities in high wavenumber modes, we set $(R_{0}/L_{Te},R_{0}/L_{Ti})=(7.0,1.0)$ instead of using $R_{0}/L_{Ti}=0.0$. We set the normalized density gradients $R_{0}/L_{na}=2.22(a=e,i)$, which is the same parameter used in the CBC caseDimits . The flux tube coordinates used here are $x=r-r_{0},y=r_{0}/q_{0}\left(q(r)\theta-\zeta\right)$, and $z=\theta$ in low-$\beta$, large aspect ratio, axisymmetric tokamak plasmas with circular, concentric flux surfaces, where $r$, $\theta$, and $\zeta$ denote the minor radius, poloidal angle, and toroidal angle, respectively. The subscript 0 signifies the parameters in the center of the flux tube.
@@
The perpendicular wavenumber is expressed as $\bm{k}=\left(k_{x}+\hat{s}zk_{y}\right)\bm{e}_{r}+k_{y}\bm{e}_{\theta}$. Here, $k_{x}$ and $k_{y}$ represent wavenumbers along the directions of $\nabla x$ and $\nabla y$, respectively, while $\bm{e}_{r}$ and $\bm{e}_{\theta}$ are unit vectors aligned with $\nabla r$ and $\nabla\theta$, respectively. The ion beta value is set to $\beta_{i}=4\pi n_{i}T_{i}/B_{ax}^{2}=1\times 10^{-4}$, ensuring the validity of the electrostatic approximation. Here, $B_{ax}$ denotes the magnetic field strength at the magnetic axis. In these simulations, $B_{\parallel\bm{k}}$ is neglected, whereas $A_{\parallel\bm{k}}$ is retained to prevent numerical difficulties arising from very rapid electrostatic waves known as the $\omega_{H}$ mode Lee . The Lenard-Bernstein collision operator Lenard is employed due to its computational efficiency compared to more sophisticated collision models. Nevertheless, we anticipate that the choice of collision model does not significantly affect the results of this study, given that the normalized collision frequency is set to $\nu^{*}_{ii}\equiv R_{0}q\nu_{ii}/(\epsilon_{r}^{3/2}v_{ti})=0.068\ (R_{0}\nu_{ii}/v_{ti}=3.7\times 10^{-3})$, which is considerably smaller than the growth rates of the instabilities examined here. The thermal velocity for particle species $a$ is defined as $v_{ta}\equiv\sqrt{T_{a}/m_{a}}$.
@@
Figure 1 shows linear growth rates and real frequencies obtained by linear simulations varying the temperature gradients. The sign of real frequencies reverses around $(R_{0}/L_{Te},R_{0}/L_{Ti})=(5.0,3.0)$, indicating that the most unstable mode transitions between ITG and TEM. For the case of $(R_{0}/L_{Te},R_{0}/L_{Ti})=(5.0,3.0)$, the most unstable mode transitions between ITG and TEM around $k_{y}\rho_{ti}=0.30$. In this region, ITG and TEM modes are comparably unstable, and therefore the solution of the initial-value problem does not converge to a single eigenmode of the form $\exp[(-i\omega_{r}+\gamma)t]$. The $k_{y}\rho_{ti}=0.30$ mode corresponds to this transition boundary and is not plotted in Fig. 1. Nonlinear simulations are performed using these plasma parameters to investigate the direction, magnitude, underlying physical mechanisms, and predictability of quasilinear models for turbulent energy exchange in ITG-TEM turbulence. Although the traditional twist-and-shift flux-tube model is known to affect the nonlinear simulation results due to the self-interaction Ball , this issue does not significantly alter our analysis and conclusions.